OpenStudy (anonymous):

Find tan2A when cosA = -3/5 and π/2

OpenStudy (dumbcow):

use right triangle to find sinA and tanA  sinA = 4/5 tanA = -4/3 now use double angle identity to find tan(2A) \[\tan (2A) = \frac{2\tan A}{1-\tan^{2} A}\] plug in -4/3 for tanA